Use parallelogram ABCD to solve the following problem. m < BAD = 4x+6 and m < BCD = 6x-8. Solve for x and the missing angles.
x =
m < A =
m < B=

Answer by ikleyn(52817)

.

In any parallelogram, the opposite angles are congruent.


So, you can write this equation


    m (< A) = m (< C)    or

    4x + 6 = 6x - 8.


Simplify and find x

     6 + 8 = 6x - 4x

       14  =    2x

        x  =     14/2 = 7.


Now calculate the angles  m (< A) = 4x + 6 = 4*7 + 6 = 34  degrees

                          m (< B) = 180 - 34 = 146  degrees.

Solved, answered and explained.